### v5

# A Survey on Diffusion Models for Time Series and Spatio-Temporal Data

Abstract.
Diffusion models have been widely used in time series and spatio-temporal data, enhancing generative, inferential, and downstream capabilities. These models are applied across diverse fields such as healthcare, recommendation, climate, energy, audio, and traffic. By separating applications for time series and spatio-temporal data, we offer a structured perspective on model category, task type, data modality, and practical application domain. This study aims to provide a solid foundation for researchers and practitioners, inspiring future innovations that tackle traditional challenges and foster novel solutions in diffusion model-based data mining tasks and applications. For more detailed information, we have open-sourced a repository [^1] https://github.com/yyysjz1997/Awesome-TimeSeries-SpatioTemporal-Diffusion-Model.
Diffusion Models, Time Series, Spatio-Temporal Data, Generative Model, Temporal Data
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ACM Comput. Surv.; 978-1-4503-XXXX-X/18/06ccs: General and reference Surveys and overviewsccs: Information systems Data mining
1. Introduction
Diffusion models represent a family of probabilistic generative models that undergo optimization through a two-step process involving the injection and subsequent removal of noise across a set of training samples. This process comprises a forward phase, referred to as diffusion, and a reverse phase, known as denoising. By training the model to remove the noise added during the diffusion process, the model learns to generate effective data samples during inference that align closely with the distribution of the training data (Ho et al., 2020; Chang et al., 2023a).
Figure 1. Trends in the cumulative number of papers related to diffusion models for time series and spatio-temporal data.
In recent years, diffusion models have risen to prominence and significantly influenced various domains, including computer vision (CV) (Ho et al., 2020; Dhariwal and Nichol, 2021; Ho et al., 2022; Saharia et al., 2022a, c), natural language processing (NLP) (Austin et al., 2021; Li et al., 2022b; Gong et al., 2022; Lovelace et al., 2024), and general multimodal learning (Ramesh et al., 2022; Saharia et al., 2022b; Yang et al., 2023d; Zhang et al., 2024). This challenges the long-time supremacy of generative adversarial networks (GANs) (Goodfellow et al., 2014; Yang et al., 2023e). Within these areas, diffusion models have demonstrated remarkable capabilities in applications such as text-to-image (Saharia et al., 2022b; Zhang et al., 2023c), instance segmentation (Amit et al., 2021; Wolleb et al., 2022b), 3D shape generation (Vahdat et al., 2022; Nam et al., 2022), molecule design (Xu et al., 2022b; Guo et al., 2023; Abramson et al., 2024), and audio generation (Huang et al., 2022; Kong et al., 2020). Remarkably, diffusion models have also gained popularity as a non-autoregressive alternative for tasks conventionally dominated by autoregressive methods (Chang et al., 2023a). Recently, the introduction of OpenAI Sora (OpenAI Sora, [n. d.]) marked the advent of diffusion models in modeling the physical world embedded within the spacetime continuum, highlighting their critical importance. In addition, AlphaFold 3 (Abramson et al., 2024) proposed by Google DeepMind uses diffusion models to generate 3D atomic coordinates and predict biomolecular structures like proteins, DNA, and RNA.
Temporal data, which primarily includes time series and spatio-temporal data, encapsulates the dynamics of the vast majority of real-world systems (Jin et al., 2023c). These forms of temporal data have been extensively studied and are recognized as crucial for numerous applications (Jin et al., 2023a; Harutyunyan et al., 2019; Jin et al., 2023b). However, deriving universal dynamic laws in the physical world from various data modalities remains a significant challenge within the field. Recently, the area of time series and spatio-temporal modeling has experienced a substantial shift from sensory intelligence towards general intelligence (Jin et al., 2024). This shift is characterized by the emergence of unified foundation models (FMs) that possess versatile temporal data analytical capabilities (Jin et al., 2023c, 2024), challenging the supremacy of domain-specific models. Diffusion models have achieved state-of-the-art results in many modalities, including images, speech, and video (Khanna et al., 2023). Benefiting from the vast and diverse available data in these fields, diffusion models often serve as generative FMs alongside large language models (LLMs) or other foundation models, facilitating rapid development in these areas (Rombach et al., 2022; Ho et al., 2022). In recent years, there has also been an increasing number of diffusion models crafted for modeling time series and spatio-temporal data (Figure 1 ). Also, we have become aware of an increasing number of attempts to use diffusion models for temporal modeling (see Table 1 ). Observing the success of diffusion models, an intriguing question arises: what kind of sparks will emerge from the intersection of time series/spatio-temporal data analysis and diffusion models?
Figure 2. An overview of diffusion models for time series and spatio-temporal data analysis. In the diffusion process, $x_{k}$ and $x_{k-1}$ denote the results after adding noise at step $k$ and k1k-1, respectively. This process can be represented by the size of the controlling steps $\beta_{k}\in(0,1)$, the identity matrix $\mathbf{I}$, and a Gaussian distribution $\mathcal{N}(x;\mu,\sigma)$ of $x$ with the mean $\mu$ and the covariance $\sigma$. During the denoising process, the model attempts to iteratively learn the data distribution by modelling the distribution ${p}_{\theta}(x_{k-1}|x_{k})$. The functions $\mu_{\theta}(\cdot)$ and variance $\sigma_{\theta}(\cdot)$ are the model learnable parameters.
Time series and spatio-temporal data analysis fundamentally rely on a profound understanding of their inherent temporal dynamics, wherein primary tasks predominantly focus on the generative capabilities of backbone models, such as forecasting (Rasul et al., 2021; Li et al., 2022a; Feng et al., 2024), imputation (Tashiro et al., 2021; Wang et al., 2023d; Yun et al., 2023; Wang et al., 2025a) and generation (Narasimhan et al., 2024; Zhang et al., 2024). These analyses center on generating temporal data samples for specific purposes in conditional or unconditional manners. Having witnessed the recent development of time series and spatio-temporal foundation models (Jin et al., 2023c; Liang et al., 2024), whether built upon LLMs or trained from scratch, their success can be attributed to the ability to estimate the distribution of training samples where effective data representations can be drawn. In this regard, diffusion models emerge as a powerful generative framework that enables (1) the modeling of complex patterns within temporal data and (2) the support of a wide range of downstream tasks, as depicted in Figure 2 .
To generate valid data samples for specific tasks, time series and spatio-temporal diffusion models usually operate in an unconditional manner without the need for supervision signals. Given the partially-observed nature of real-world applications (Wang et al., 2024a), conditional diffusion models have emerged. They leverage data labels (e.g., instructions, metadata, or exogenous variables) to regulate the generation process, thereby enabling effective cross-modal prompting that leads to more tailored and improved outcomes (Narasimhan et al., 2024). We present a roadmap in Figure 3 . By training on large-scale temporal data, diffusion models effectively fill the gap in time series/spatio-temporal data generation and exhibit significant potential in solving the puzzle of next-generation, LLM-empowered temporal data-centric agents (Wu et al., 2023; Jin et al., 2024).
Figure 3. Representative diffusion models for time series and spatio-temporal data in recent years.
Despite the promising prospects and rapid advancement of diffusion models in handling time series and spatio-temporal data, there has been a conspicuous lack of systematic analysis of this model family in the existing literature. This article aims to bridge this gap by providing a forward-looking review that elucidates both the ’why’ and the ’how’, detailing the reasons diffusion models are suited for these data modalities and unveiling the mechanisms through which they confer advantages. In this survey, we offer a detailed categorization, engage in thorough reviews, and identify burgeoning trends within this rapidly evolving landscape. Our main contributions are summarized as follows.
  • Comprehensive and up-to-date review. We present a comprehensive, up-to-date, and forward-looking review of diffusion models for time series and spatio-temporal data. Our survey highlights the suitability of diffusion models for these data modalities and discusses the benefits they confer. By covering both a broad spectrum of the field and the specifics of individual methods, we furnish readers with a deep insight into this subject area.
  • Unified and structured categorization. We introduce a clear and organized framework for categorizing the existing literature into two main types: unconditional and conditional diffusion models, focusing on time series and spatio-temporal data that span both predictive and generative tasks. This categorization offers the reader a coherent roadmap of the topic from multiple perspectives.
  • Insights into emerging advances. We discuss cutting-edge techniques in both unconditional and conditional diffusion models, focusing on time series and spatio-temporal data. Our coverage includes the latest techniques and emerging trends such as multimodal conditional generation.
  • Summary of challenges and future directions. We identify key challenges faced in the current research landscape and highlight several promising directions for future exploration.
The remainder of this paper is structured as follows: section 2 provides a comprehensive background on diffusion models, detailing their development, theoretical foundations, and various implementations. section 3 presents a structured overview and categorization of diffusion models applied to time series and spatio-temporal data, setting the stage for a deeper exploration of model perspectives in section 4 , which discusses both standard and advanced diffusion models. section 5 focuses on task perspectives, examining how diffusion models tackle forecasting, generation, imputation, anomaly detection, and more. section 6 discusses data perspectives, highlighting challenges and solutions specific to time series and spatio-temporal data. section 7 explores the application of diffusion models across various domains, such as healthcare, traffic, and energy, demonstrating their broad utility. Finally, section 8 concludes the paper with an outlook on future opportunities and summarizing remarks.
2. Background
This paper primarily reviews recent improvements in using diffusion models to solve time series and spatio-temporal data challenges. In this section, we will first define time series and spatio-temporal data, as well as their corresponding tasks in various fields. Then, we will introduce the history of the diffusion model and its advantages. Finally, some different kinds of diffusion models and their variants based on theoretical formula derivations and comparisons with other generative models will be presented.
2.1. Overview of Time Series and Spatio-Temporal Data
Temporal data, particularly time series and spatio-temporal data, are important data structures for a wide range of real-world applications (Jin et al., 2023c). A time series is defined as a sequential arrangement of data points, categorized by their temporal order. These sequences may be univariate, involving a single variable over time, or multivariate, incorporating multiple variables. For instance, daily air quality measurements in a city constitute a univariate time series, while combining daily temperature and humidity readings generates a multivariate series. In our discussions, we employ bold uppercase letters (e.g., $\mathbf{X}$) to represent matrices, bold lowercase (e.g., $\mathbf{x}$) for vectors, calligraphic uppercase (e.g., $\mathcal{X}$) for sets, and standard lowercase (e.g., $x$) for scalars.
We base our formal definitions of time series on those provided in (Jin et al., 2023c, a). Specifically, a univariate time series (Figure 4 (a)), denoted as $\mathbf{x}=(x_{1},x_{2},\ldots,x_{T})\in\mathbb{R}^{T}$, consists of a sequence of $T$ data points arranged chronologically, where each $x_{t}\in\mathbb{R}$ represents the series’ value at time $t$. Conversely, a multivariate time series (Figure 4 (b)), represented by $\mathbf{X}=(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T})\in\mathbb{R}^{T\times D}$, encompasses a sequence of $T$ data points also in temporal sequence but across $D$ feature channels, with $\mathbf{x}_{t}\in\mathbb{R}^{D}(1\leq t\leq T)$ indicating the series’ values at time $t$ across $D$ different channels. For a comprehensive exploration of time series, we direct the reader to (Jin et al., 2023a).
Spatio-temporal data, in contrast, encompasses sequences of data points characterized by both their temporal and spatial dimensions. This type of data integrates the aspect of time, as seen in time series, with the additional dimension of space, capturing the complex dynamics of phenomena as they unfold over time and across different locations. From this perspective, multivariate time series may also be considered as a form of spatio-temporal data. Such data is instrumental in fields ranging from geography and meteorology to urban planning and environmental monitoring, where understanding the interplay between spatial patterns and temporal evolution is crucial.
Figure 4. Illustrations of time series and spatio-temporal data.
In practical applications, spatio-temporal data refers to a collection of observations where each data point is defined by its position in space and time, encapsulating a diverse range of data structures such as graphs, trajectories, and even videos, as noted in (Jin et al., 2023c).
For instance, a spatio-temporal graph (Figure 4 (c)) representing urban traffic flow (Sun et al., 2020; Liang et al., 2021a) over time can be understood as spatio-temporal data where each node represents a specific location with specific attributes and edges are weighted by traffic volume, which changes over time. Likewise, trajectory data (Liang et al., 2021b; Chen et al., 2024c) (Figure 4 (d)) captures the movement of objects through space over time, including their paths, speeds, and changes in direction. Such data is vital for applications in transportation studies, wildlife tracking, and mobile network optimization, where analyzing the patterns of movement and predicting future locations based on historical data are of paramount importance. On this basis, these and other spatio-temporal constructs can be systematically characterized and analyzed.
Building on the definitions provided, we now proceed to succinctly introduce the representative tasks associated with each data category (Jin et al., 2023c).
  • Time Series Analysis. The analysis of time series with diffusion models encompasses four primary tasks: forecasting, generation, anomaly detection, and imputation. Forecasting focuses on predicting future values within a time series, which can be subdivided into short-term and long-term forecasts based on the temporal scope of the predictions. Generation involves creating new time series based on the statistical properties of a given dataset, serving as a way to simulate possible scenarios or enhance data diversity for training models. Anomaly detection, a specialized form of classification, aims to distinguish atypical series from normal ones. Imputation addresses the challenge of filling in missing values within a series, which is crucial for maintaining the integrity and utility of time series.
  • Spatio-Temporal Data Analysis. Spatio-temporal data analysis, while encompassing tasks similar to those in time series analysis, often applies these methodologies within specific application scenarios. For instance, forecasting may focus on traffic flow (Wen et al., 2023) or air quality (Chen et al., 2024b), utilizing historical data patterns to predict future conditions. Generation tasks might involve creating synthetic trajectories, offering privacy-compliant alternatives to original datasets by replacing sensitive information with generated, anonymized data (Zhu et al., 2024a, b). Anomaly detection becomes particularly crucial in scenarios such as vehicle trajectory analysis, where deviations from generated normative patterns may indicate unusual or suspicious behaviors (Li et al., 2024a). Additionally, spatio-temporal imputation plays a vital role in addressing missing values in multivariate time series (Yun et al., 2023; Du et al., 2024), ensuring comprehensive and accurate datasets for further analysis. More discussion is given in section 5 .
2.2. Why Diffusion Model and Its History
Diffusion models are a class of probability-based generative models. They are called after the mathematical process of diffusion, which is commonly used to describe phenomena such as particle movement in a gas or liquid (Sohl-Dickstein et al., 2015). In detail, the concept of diffusion models first appeared in statistical physics, used to describe the process of particles moving from areas of high concentration to areas of low concentration (Jarzynski, 1997). Early diffusion models were primarily concerned with accurately simulating the random diffusion behavior in the generation process.
One of the key breakthroughs in diffusion models occurred in 2015, when researchers proposed a method that combines variational inference to effectively train these models (Sohl-Dickstein et al., 2015). Since then, the field has experienced rapid development, especially in the area of high-resolution image generation (Rombach et al., 2022). Since 2020, diffusion models have begun to show their potential in more fields, such as text2image, music generation, and speech synthesis (Saharia et al., 2022b; Wang, 2023; Richter et al., 2023). These advances are due to the optimization of model structures, improvements in training methods, and increased computational resources.
In addition to the field of application, in terms of theory, researchers began to explore how to generate data with specific features by controlling the reverse process. This study direction finally resulted in the development of diffusion models capable of producing high-quality, multidimensional data. In updated computer algorithms, it can be represented as the process of gradually modifying the data distribution, progressively injecting noise until it matches the target distribution, in order to obtain high-quality and realistic synthetic data samples (Ho et al., 2020).
Recently, more and more researchers and engineers are now focusing their perspectives on the diffusion model, and it has become one of the first choices for generating models (Yang et al., 2023e). Diffusion models excel at generating high-quality, complex sequences, including time series and spatio-temporal data, with detailed coherence by gradually removing noise. They offer strong control over generations, allowing fine-tuning based on the conditions (Croitoru et al., 2023). These models are flexible and adaptable across various data types and modalities, robust against errors with a gradual noise reduction mechanism, and capable of exploring data diversity for creative outputs. Moreover, they can be integrated with other model types, like autoencoders, to enhance generation quality and control (Yang et al., 2023e).
2.3. Typical Diffusion Models
Typically, the training process includes two steps: the forward process (diffusion) and the reverse process (denoising). Diffusion models start with a noise distribution, which is gradually altered through a series of steps. In the forward process, the model incrementally adds noise to the original data over multiple steps until the data turns into pure random noise. This process is usually Markovian, meaning that each step depends only on the preceding one. Then the reverse process takes place, involving learning to remove the noise from the data, essentially reversing the forward process. By training the model to remove the noise added during the diffusion process, the model learns to generate samples from the same distribution as the training data. The entire training process involves optimizing the model to denoise effectively. This is typically done using a loss function that encourages the model to produce samples that are close to the true data distribution (Nichol and Dhariwal, 2021).
The currently common frameworks for diffusion models include denoised diffusion probabilistic models (DDPMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020), score-based stochastic differential equations (Score SDEs) (Song et al., 2020b; Song and Ermon, 2019), conditional diffusion models (Rombach et al., 2022; Dhariwal and Nichol, 2021; Ho and Salimans, 2022), etc. Following, we will introduce the subclasses of the diffusion model through theoretical formula derivation.
2.3.1. Denoised Diffusion Probabilistic Models (DDPMs)
Denoised diffusion probabilistic models are built around a well-defined probabilistic process via dual Markov chains that consist of two parts: a diffusion (or forward) process that gradually transforms the data into noise with pre-determined noise, such as Gaussian noise, and a denoising (or reverse) process that attempts to recover the original data by deep neural networks.
Forward (Diffusion) Process. Given a data distribution $q(\mathbf{x})$ and sample an initial clean data $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$ from it. The subsequent forward diffusion process incrementally adulterates the initial data distribution by superimposing Gaussian noise, and finally progresses towards convergence with the standard Gaussian distribution. In the diffusion process up to step $K$, a sequence of distributed latent data $\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{K}$ materializes. The diffusion process can be defined as a Markov chain transforms $\mathbf{x}_{k-1}$ to $\mathbf{x}_{k}$ with a diffusion transition kernel:
(1) $q(\mathbf{x}_{k}|\mathbf{x}_{k-1}):=\mathcal{N}(\mathbf{x}_{k};\sqrt{1-\beta_{k}}\mathbf{x}_{k-1},\beta_{k}\mathbf{I}),$
for $\forall{k}\in\{1,\cdots,K\}$ with the size of the controlling steps $\beta_{k}\in(0,1)$, the identity matrix $\mathbf{I}$, and a Gaussian distribution $\mathcal{N}(\mathbf{x};\mu,\sigma)$ of $\mathbf{x}$ with the mean $\mu$ and the covariance $\sigma$. According to the properties of the Gaussian kernel, it is feasible to get $\mathbf{x}_{k}$ directly from $\mathbf{x}_{0}$ by equation 1, and collect noise samples straight from the original input $\mathbf{x}_{0}$ for any step, that is,
(2) $q(\mathbf{x}_{k}|\mathbf{x}_{0}):=\prod_{k=1}^{K}{q}(\mathbf{x}_{k}|\mathbf{x}_{k-1}):=\mathcal{N}(\mathbf{x}_{k};\sqrt{\bar{\alpha_{k}}}\mathbf{x}_{0},\sqrt{1-\bar{\alpha_{k}}}\mathbf{I}),$
with $\alpha_{k}:=1-\beta_{k}$ and $\bar{\alpha_{k}}:=\prod_{i=1}^{K}\alpha_{i}$. Therefore, $\mathbf{x}_{k}=\sqrt{\bar{\alpha_{k}}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha_{k}}}\epsilon$ with Gaussian noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. Typically, it is designed $\bar{\alpha_{k}}\approx 0$, s.t., $q(\mathbf{x}_{k}):=\int q(\mathbf{x}_{K}|\mathbf{x}_{0})q(\mathbf{x}_{0})\mathbf{d}\mathbf{x}_{0}\approx\mathcal{N}(\mathbf{x}_{k};\mathbf{0},\mathbf{I})$, i.e., the backward chain can begin with any Gaussian noise. Overall, the forward process gradually injects noise into the data until all structures have disappeared.
Reverse (Denoising) Process. The reverse process performs the denoising task at each step with a series of Markov chains until the damaged original data is reconstructed. Specifically, the series of reverse Markov chains start with a distribution $p(\mathbf{x}_{K})=\mathcal{N}(\mathbf{x}_{K};\mathbf{0},\mathbf{I})$ and a learnable kernel $p_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k})$ to generate $p_{\theta}(\mathbf{x}_{0})$. The learnable Gaussian transition kernels $p_{\theta}$ can be represented as:
(3) ${p}_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k}):=\mathcal{N}(\mathbf{x}_{k-1};\mu_{\theta}(\mathbf{x}_{k},{k}),\sigma_{\theta}(\mathbf{x}_{k},{k})\mathbf{I}),$
where the mean $\mu_{\theta}(\cdot)$ and variance $\sigma_{\theta}(\cdot)$ are the model learnable parameters. The model tries to learn the data distribution by the model distribution ${p}_{\theta}(\mathbf{x}_{0})$ during the reverse denoising process.
Training. In order to approximate the real data distribution, the diffusion model is trained to minimize variational constraints on the negative log-likelihood (NLL):
(4) $\mathbb{E}\left[-\log p_{\theta}\left(\mathrm{x}_{0}\right)\right]$ $\leq\mathbb{E}_{q}\left[-\log\frac{p_{\theta}\left(\mathrm{x}_{0:K}\right)}{q\left(\mathrm{x}_{1:K}\mid\mathrm{x}_{0}\right)}\right]=\mathbb{E}_{q}\left[-\log p\left(\mathrm{x}_{K}\right)-\sum_{k\geq 1}\log\frac{p_{\theta}\left(\mathrm{x}_{k-1}\mid\mathrm{x}_{k}\right)}{q\left(\mathrm{x}_{k}\mid\mathrm{x}_{k-1}\right)}\right]=:L.$
It is equivalent to Kullback–Leibler divergence (KL divergence) format as mentioned in the DDPM paper (Ho et al., 2020) with three parts: the prior loss $L_{K}$, the divergence of the forwarding step and the corresponding reversing step $L_{k-1}$, and the reconstruction loss $L_{0}$:
(5) $L$= $\mathbb{E}_{q}[\underbrace{D_{\mathrm{KL}}\left(q\left(\mathrm{x}_{K}\mid\mathrm{x}_{0}\right)\|p\left(\mathrm{x}_{K}\right)\right)}_{L_{K}}+\sum_{k>1}\underbrace{D_{\mathrm{KL}}\left(q\left(\mathrm{x}_{k-1}\mid\mathrm{x}_{k},\mathrm{x}_{0}\right)\|p_{\theta}\left(\mathrm{x}_{k-1}\mid\mathrm{x}_{k}\right)\right)}_{L_{k-1}}-\underbrace{\log p_{\theta}\left(\mathrm{x}_{0}\mid\mathrm{x}_{1}\right)}_{L_{0}}].$
Especially, to minimize the NLL, we can only train the divergence loss between two steps $L_{k-1}$, and using Baye’s rule to parameterize the posterior $q\left(\mathrm{x}_{k-1}\mid\mathrm{x}_{k},\mathrm{x}_{0}\right)$, that is:
(6) $q\left(\mathrm{x}_{k-1}\mid\mathrm{x}_{k},\mathrm{x}_{0}\right)=\mathcal{N}\left(\mathrm{x}_{k-1};\tilde{\boldsymbol{\mu}}_{k}\left(\mathrm{x}_{k},\mathrm{x}_{0}\right),\tilde{\beta}_{k}{I}\right),$
(7) $\tilde{\mu}_{t}(\mathbf{x}_{k},\mathbf{x}_{0}):=\frac{\sqrt{\bar{\alpha}_{k-1}}\beta_{k}}{1-\bar{\alpha}_{k}}\mathbf{x}_{0}+\frac{\sqrt{\alpha_{k}}(1-\bar{\alpha}_{k-1})}{1-\bar{\alpha}_{k}}\mathbf{x}_{k}\hskip 2.84526pt,$
(8) $\tilde{\beta}_{k}:=\frac{1-\bar{\alpha}_{k-1}}{1-\bar{\alpha}_{k}}\beta_{k}.$
where $\alpha_{k}$ is $1-\beta_{k}$ and $\bar{\alpha}_{k}$ indicates $\prod_{k=1}^{K}\alpha_{k}$. $L_{k-1}$ can be equated to the expected value of the $\ell_{2}$-loss between the two mean coefficients:
(9) $L_{k-1}=\mathbb{E}_{q}\left[\frac{1}{2\sigma_{k}^{2}}\left\|\tilde{\boldsymbol{\mu}}_{k}\left(\mathrm{x}_{k},\mathrm{x}_{0}\right)-\boldsymbol{\mu}_{\theta}\left(\mathrm{x}_{k},k\right)\right\|^{2}\right]+C.$
Ho et al. (Ho et al., 2020) emphasize that, rather than parameterizing the mean $\mu_{\theta}(\mathrm{x}_{k},k)$, predicting the noise vector at each time step in the forward process by parameterizing $\epsilon_{\theta}(\mathrm{x}_{k},k)$ for simplification:
(10) $\mathbb{E}_{k\sim\mathcal{U}(1,K),\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\bigg[{\lambda(k)\left\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{k},k)\right\|^{2}}\bigg],$
where $\lambda(k)=\frac{{\beta_{k}}^{2}}{2{\sigma_{K}}^{2}\alpha_{k}(1-\bar{\alpha}_{k})}$ is a weight that changes the noise scale and $\epsilon_{\theta}$ is a model for Gaussian noise prediction. After training using the above loss function, $\epsilon_{\theta}$ will be used in the reverse process of ancestor sampling.
Inference (Sampling). Given the noisy data $\mathrm{x}_{K}$ and starting with step $K$ denoising, the final time series is generated through the equation:
(11) $p_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k})$ $=\mathcal{N}(\mathbf{x}_{k-1};\mu_{\theta}(\mathbf{x}_{k},k),\sigma_{\theta}^{2}(\mathbf{x}_{k},k)\mathbf{I})\sim\frac{1}{\sqrt{\alpha_{k}}}(\mathbf{x}_{k}-\frac{\beta_{k}}{\sqrt{1-\overline{\alpha_{k}}}}\epsilon_{\theta}(\mathbf{x}_{k},k))+\sigma_{\theta}(\mathbf{x}_{k},k)z,$
where $z\sim\mathcal{N}(0,\mathbf{I})$, also $\beta_{k}\approx\sigma_{\theta}^{2}(\mathbf{x}_{k},k)$ in practice.
2.3.2. Score SDE Formulation
DDPM achieved a set of discrete steps in the forward processing, so it has some limitations about training designs. Score SDE further generalizes DDPM’s discrete system to a continuous framework based on the stochastical differential equation (Song et al., 2020b). Here we use $T$ instead of the step size $k$ in DDPM.
Forward Process. The corresponding continuous diffusion process can be represented using Itô SDE (Ito et al., 1951), including a mean shift and a Brownian motion (standard Wiener process) as follows:
(12) $\mathrm{d}\mathbf{x}=\mathbf{f}(\mathbf{x},t)\mathrm{d}t+g(t)\mathrm{d}\mathbf{w},t\in[0,T],$
where $\mathbf{f}(\cdot,t)$ represents the drift coefficient for the stochastic process $\mathbf{x}(t)$, and $g(\cdot)$ is the diffusion coefficient linked to the Brownian motion $\mathbf{w}$.
Reverse Process. The new samples can be synthesized from the known prior distribution $p_{T}$ by solving the reverse-time SDE (Anderson, 1982):
(13) $\mathrm{d}\mathbf{x}=\left[\mathbf{f}(\mathbf{x},t)-g^{2}(t)\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right]\mathrm{d}t+g(t)\mathrm{d}\bar{\mathbf{w}},$
where $\bar{\mathbf{w}}$ is a Brownian motion with reversed time flows (Vincent, 2011). The solution to the reverse-time SDE is approximated by a time-dependent neural network $s_{\theta}(\mathbf{x},t)$ to a score function $\nabla_{\mathbf{x}}\log p_{t}{(\mathbf{x})}$. The solution of the forward SDE equation is that:
(14) ${L}=\mathbb{E}_{t}$ $\{\lambda(t)\mathbb{E}_{\mathbf{x}_{0}}\mathbb{E}_{q(\mathbf{x}_{t}|\mathbf{x}_{0})}[\|{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t}t)-\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}|\mathbf{x}_{0})\|_{2}^{2}]\},$
where $\mathbf{x}_{0}$ is sampled from the distribution $p_{0}$ and $\lambda(t)$ is the positive weighting function. This method circumvents the direct approximation of the computationally infeasible score function by estimating the transition probability that follows a Gaussian distribution during the forward diffusion process (Song et al., 2020b).
Besides, there are some simplified explanations of reverse-time SDE solvers. Using these techniques, samples can be generated after training using various methods. Euler-Maruyama (EM) Method (Mao, 2015): Solves the reverse-time SDE through a simple discretization technique, replacing $\mathrm{d}\mathbf{x}$ with $\triangle\mathbf{t}$ and $\mathrm{d}\bar{\mathbf{w}}$ with Gaussian noise $z$. Prediction-Correction (PC) Method: Operates in a sequential manner, alternating between predictor and corrector steps. The predictor can use any numerical solver, such as the EM method, for the reverse-time SDE, while the corrector can be any score-based Markov Chain Monte Carlo (MCMC) method. Probability Flow ODE Method (Song et al., 2020b): Reformulates the forward SDE into an ODE that maintains the same marginal probability density $p_{t}$ as the SDE. Sampling by solving this reverse-time ODE is equivalent to solving the time-reversed SDE. There are also some advanced ODE solvers to speed up the process (Liu et al., 2022c; Lu et al., 2022b).
2.3.3. Conditional Diffusion Models
In the previous sections, we discussed the DDPM and Score SDE from an unconditional view. They generate data without any explicit conditions or guidance. The model learns to produce outputs from the learned distribution of the input data. The general diffusion models are capable of generating data samples not just from an unconditional distribution $p_{0}$, but also from a conditional distribution $p_{0}(\mathbf{x}|c)$ when given a condition $c$. This condition could be class labels or features related to the input data $\mathbf{x}$ (Rombach et al., 2022). During training, the score network $s_{\theta}(\mathbf{x},t,c)$ takes condition $c$ as an input. Additionally, there are specific sampling algorithms designed for conditional generation, such as label-based conditions (Dhariwal and Nichol, 2021), label-free conditions (Ho and Salimans, 2022), and further distillation-based guidance (Dhariwal and Nichol, 2021), self guidance (Epstein et al., 2024; Yang et al., 2022), textual-based guidance (Le et al., 2024; Gong et al., 2022), graph-based guidance (Schneuing et al., 2022), physical-based guidance (Yuan et al., 2023b), task-based guidance (An et al., 2023), etc. These conditional mechanisms are more conducive to the generation of application-specific fields by using the control of other information to generate results (Yang et al., 2023e).
In detail, sampling under labels and classifiers’ conditions involves using gradient guidance at each step, which typically requires an additional classifier with encoder architecture (e.g., U-Net (Ronneberger et al., 2015) and Transformer (Vaswani et al., 2017)) to generate condition gradients for specific labels (Dhariwal and Nichol, 2021). These labels are flexible and can be textual or categorical, binary, or based on extracted features (Dhariwal and Nichol, 2021; Lu et al., 2022b; Nichol et al., 2021; Meng et al., 2023; Hu et al., 2022; Wolleb et al., 2022b; Chen et al., 2023b; Baranchuk et al., 2021; Yang et al., 2024b; Zuo et al., 2023). Correspondingly, sampling under unlabeled conditions relies solely on self-information for guidance (Epstein et al., 2024; Chao et al., 2022; Kollovieh et al., 2024). Compared to the high accuracy of the labeled conditional diffusion model, the unlabeled one has advantages in generating innovative and diverse data. Therefore, unlabelled models are more suitable for exploratory and creative application scenarios (Chung et al., 2022; Choi et al., 2021). Furthermore, there are also more conditional methods being proposed, which use information about the data itself (Kollovieh et al., 2024), other modalities (Nichol et al., 2021; Liu et al., 2023a), other representations (Yuan et al., 2023b), and other knowledge (Shao et al., 2024) as conditions to guide the diffusion model for generation. Currently, the condition-based diffusion model is also the most common method in various application scenarios due to its highly specific and controlled outputs (Bansal et al., 2023; Zhang et al., 2023c).
2.3.4. Improvements with Diffusion Model and Its Variants
While diffusion models have yielded satisfactory results in various tasks, practical applications reveal limitations such as slow iterative sampling and computational complexity due to high-dimensional input. There are also concerns regarding generalization performance and integration with other generative models. Next, we will explore the variants of diffusion models from efficiency and performance improvement, highlighting the modifications and optimizations they bring in comparison to the original diffusion models.
Efficiency Enhancement. (1) Forward Processes Improvement: Several innovations aim to enhance model efficiency and robustness by using physical phenomena. PFGM (Xu et al., 2022a) and its extension PFGM++ (Xu et al., 2023b) and (Liu et al., 2023d) use electric field dynamics and augmented dimensions for improved performance. Cold Diffusion uses image transformations as a forward process (Bansal et al., 2022), and improved Gaussian perturbation kernel methods were presented (Karras et al., 2022). (2) Reverse Processes Improvement: Methods to reduce generation steps or use lightweight models include training-free sampling methods that optimize the trajectory from noise to data distributions, such as DDIM (Song et al., 2020a), gDDIM (Zhang et al., 2022), PNDM (Liu et al., 2022c), EDM (Karras et al., 2022), DEIS (Zhang and Chen, 2022), and DPM-Solver (Lu et al., 2022b). SDE-based methods use techniques like restart sampling (Jolicoeur-Martineau et al., 2021; Xu et al., 2023a). Knowledge distillation is also used by transferring insights from larger models to simpler ones (Berthelot et al., 2023; Song et al., 2023). (3) Integration with Other Models: Combining diffusion models with VAE (Lyu et al., 2022; Pandey et al., 2022) and GAN (Zheng et al., 2022; Watson et al., 2021) improves efficiency. Using latent space as input can reduce dimensionality and enhance efficiency (Pandey et al., 2022; Rombach et al., 2022; Vahdat et al., 2021; Zhang et al., 2023b). (4) Scheduler Functions: Optimizing the reverse process with improved schedulers such as CMS (Song et al., 2023), DDIM (Song et al., 2020a), IDDPM (Nichol and Dhariwal, 2021), DEIS (Zhang and Chen, 2022), DPM-Solvers (Lu et al., 2022b, c), Euler and Heun scheduler (Karras et al., 2022), LCM (Luo et al., 2023), RePaintScheduler (Lugmayr et al., 2022), TCD (Zheng et al., 2024), UniPC (Zhao et al., 2024a), and VQD (Gu et al., 2022) enables faster convergence and reduces iterations.
Performance Enhancement. (1) Model Architecture: Traditional DDPM uses the U-Net architecture for efficiency. Innovations include normalization (Kawar et al., 2022; Berthelot et al., 2023) and attention mechanisms with position encoding (Zhang and Chen, 2022; Rissanen et al., 2022; Kingma et al., 2021). Transformer-based architectures are increasingly used to improve performance (Vaswani et al., 2017), including ViT (Dosovitskiy et al., 2020), SwinTransformer (Liu et al., 2021a, 2022b), DiT (Peebles and Xie, 2023), and SiT (Ma et al., 2024a). (2) Training Schedules: Advanced decoding strategies involve optimizing diffusion stages and innovative projection methods. Optimization strategies adjust the diffusion steps (Zheng et al., 2022; Lyu et al., 2022), while projection techniques use various diffusion processes to increase versatility (Hoogeboom and Salimans, 2022; Daras et al., 2022). Innovations in noise adjustment (Kingma et al., 2021; Kong and Ping, 2021; Nichol and Dhariwal, 2021) and optimization methods focusing on loss and matching (Song et al., 2021; Huang et al., 2021) enhance model performance. (3) Data Bridging: To address generating arbitrary Gaussian distributions with complex distributions, research uses SDE/ODE principles. Techniques like $\alpha$-blending (Heitz et al., 2023) and Rectified Flow (Liu et al., 2022a; Albergo and Vanden-Eijnden, 2022) explore ODE creation between distributions. The Schrödinger Bridge concept offers new avenues in distribution transportation (Su et al., 2022; Liu et al., 2023e). (4) Conditional and Guidance Strategies: These strategies are crucial for directing generation and enhancing output relevance and quality in response to specific conditions, as discussed in Sec 2.3.3.
Currently, institutions like OpenAI and Stability AI have launched numerous outstanding diffusion models. For instance, Stable Diffusion generates high-quality images from textual descriptions (Rombach et al., 2022), and ControlNet enhances control over generated images (Zhang et al., 2023c). The consistency model offers state-of-the-art performance without the slow iterative process of traditional models (Song et al., 2023).
2.4. Diffusion Model vs Other Generative Models
Figure 5. Different types of generative models.
In addition to the diffusion model, there are many classical generative models. We take the most widely used variational autoencoders (VAEs), generative adversarial networks (GANs), and flow-based generative models as examples to introduce them and analyse their advantages, disadvantages and differences with the diffusion model. The workflow of these models is shown in Figure 5 .
VAE is a probabilistic generative model that encodes input data into a latent space $z$ and decodes from that latent space to generate data (Kingma and Welling, 2013), as shown in Figure 5 (a). VAEs aim to maximize the lower bound of the log-likelihood of the data, known as the Evidence Lower Bound (ELBO). GAN consists of a generator $G(z)$ and a discriminator $D(x,x^{\prime})$, as shown in Figure 5 (b). The generator tries to produce samples as close to real data $x$ as possible, while the discriminator tries to distinguish between real data $x$ and generated data $x^{\prime}$. The training of GANs involves a zero-sum game, continually optimizing both the generator and the discriminator (Goodfellow et al., 2014). Flow-based generative model (normalizing flows) transform data into a simpler distribution (e.g., Gaussian) through a sequence of reversible transformations, as shown in Figure 5 (c), ensuring precise likelihood evaluation due to the transformations being invertible (Rezende and Mohamed, 2015).
Different generative models have their unique strengths and limitations, making them suitable for specific applications. VAE is appreciated for its simplicity, stability, and clear theoretical foundation, but tends to produce lower-quality data and has limited expressiveness in its latent spaces. GAN is known for its powerful generative capabilities but is notoriously difficult to train and is prone to instability and mode collapse. Flow-based models offer precise likelihood estimation and high-quality generation but require significant computational resources and involve complex model designs. Diffusion models excel in generating high-quality, detailed output, offering flexibility and solid probabilistic foundations, but suffer from long training times and high computational costs. However, many variants and improvements of diffusion models have mitigated these issues, as described in Sec. 2.3.4.
3. Overview and Categorization
Table 1. Summary and main papers of the diffusion models for time series and spatio-temporal data modeling. Red indicates univariate time series, blue-violet is multivariate time series, and yellow shows spatio-temporal data.
Method Data Model Task Application Institute Venue Year
WaveGrad (Chen et al., 2020) Univariate DDPM Generation Audio Google ICLR 2021
DiffWave (Kong et al., 2020) Univariate DDPM Generation Audio UCSD ICLR 2021
D-Va (Koa et al., 2023) Univariate DDPM Forecasting Finance NUS CIKM 2023
DiffLoad (Wang et al., 2023b) Univariate DDPM Forecasting Electricity HKU ArXiv 2023
DiffuASR (Liu et al., 2023g) Univariate DDPM Generation Recommendation XJTU CIKM 2023
DiffRec (Wang et al., 2023c) Univariate DDPM Generation Recommendation NUS SIGIR 2023
TimeGrad (Rasul et al., 2021) Multivariate DDPM Forecasting General Zalando ICML 2021
CARD (Han et al., 2022) Multivariate DDPM Classification General UT-Austin NeurIPS 2022
BVAE (Li et al., 2022a) Multivariate DDPM Forecasting General Baidu NeurIPS 2023
SSSD (Alcaraz and Strodthoff, 2022) Multivariate DDPM Imputation General Oldenburg TMLR 2023
DA-TASWDM (Xu et al., 2023c) Multivariate DDPM Imputation Healthcare HKBU CIKM 2023
DiffEEG (Shu et al., 2023) Multivariate DDPM Forecasting Healthcare USTC ArXiv 2023
Tosato et al. (Tosato et al., 2023) Multivariate DDPM Classification Healthcare Tilburg Synapsium 2023
DiffCharge (Li et al., 2024e) Multivariate DDPM Generation Electricity HKUST ArXiv 2023
Yang et al. (Yang et al., 2023a) Multivariate DDPM Imputation&Anomaly Detection AIOps Microsoft ESEC/FSE 2023
AnoDDPM (Sui et al., 2024) Multivariate DDPM Anomaly Detection General Beihang IEEE Sens. J. 2024
DiffShape (Liu et al., 2024b) Multivariate DDPM Classification General SCUT AAAI 2024
STPP (Yuan et al., 2023a) Spatio-temporal DDPM Forecasting General Event THU KDD 2023
ERDiff (Wang et al., 2024b) Spatio-temporal DDPM Alignment General GIT NeurIPS 2023
SpecSTG (Lin et al., 2024) Spatio-temporal DDPM Forecasting Transportation USYD ArXiv 2024
DST-DDPM (Chen et al., 2024b) Spatio-temporal DDPM Forecasting Environment HKUST ENVIRON RES 2024
UTD-PTP (Tang et al., 2024) Spatio-temporal DDPM Forecasting Transportation BIT IEEE Sens. J. 2024
SGMSE (Welker et al., 2022) Univariate Score-based Generation Audio UHH Interspeech 2022
DeScoD-ECG (Li et al., 2023b) Univariate Score-based Denoising Healthcare U of A IEEE J BIOMED HEALTH 2023
ScoreGrad (Yan et al., 2021) Multivariate Score-based Forecasting General BIT ArXiv 2021
Bilo et al. (Biloš et al., 2022) Multivariate Score-based Forecasting General TUM ICML 2023
StoRM (Lemercier et al., 2023) Multivariate Score-based Denoising&Generation Audio UHH IEEE-ACM T AUDIO SPE 2023
Risk-sensitive SDE (Li et al., 2024c) Multivariate Score-based Generation General Cambridge ArXiv 2024
TimeADDM (Hu et al., 2024) Multivariate Score-based Anomaly Detection General HFUT ICASSP 2024
Sasdim (Zhang et al., 2023d) Spatio-temporal Score-based Imputation General CSU ArXiv 2023
Dyffusion (Rühling Cachay et al., 2024) Spatio-temporal Score-based Forecasting General UCSD NeurIPS 2023
Lu et al. (Lu et al., 2022a) Univariate Conditional Generation Audio CMU ICASSP 2022
PulseDiff (Jenkins et al., 2023) Univariate Conditional Imputation Healthcare IC ArXiv 2023
DR-DiffuSE (Tai et al., 2023) Univariate Conditional Generation Audio UESTC AAAI 2023
Dose (Tai et al., 2024) Univariate Conditional Generation Audio UESTC NeurIPS 2023
FUSE (Yang et al., 2024a) Univariate Conditional Generation Audio Oxford Interspeech 2024
DiffsFormer (Gao et al., 2024) Univariate Conditional Imputation&Generation Finance USTC ArXiv 2024
DreamRec (Yang et al., 2024b) Univariate Conditional Generation Recommendation USTC NeurIPS 2024
CSDI (Tashiro et al., 2021) Multivariate Conditional Imputation General Stanford NeurIPS 2021
TimeDiff (Shen and Kwok, 2023) Multivariate Conditional Forecasting General HKUST ICML 2023
DiffAD (Xiao et al., 2023) Multivariate Conditional Anomaly Detection General HENU KDD 2023
D3(Wang et al., 2024c) Multivariate Conditional Anomaly Detection General BUPT NeurIPS 2023
MIDM (Wang et al., 2023d) Multivariate Conditional Imputation General USTC KDD 2023
MEDiC (Sharma et al., 2023) Multivariate Conditional Imputation Healthcare IIT NeurIPS 2023
DiffPLF (Li et al., 2024f) Multivariate Conditional Forecasting Electricity HKUST(GZ) ArXiv 2024
ImDiffusion (Chen et al., 2023c) Multivariate Conditional Anomaly Detection General PKU VLDB 2024
DiffDA (Huang et al., 2024) Multivariate Conditional Forecasting Climate ETH Zurich ArXiv 2024
DiffSTOCK (Daiya et al., 2024) Multivariate Conditional Forecasting Finance Purdue ICASSP 2024
RF-Diffusion (Chi et al., 2024) Multivariate Conditional Generation Network THU MobiCom 2024
DiffSTG (Wen et al., 2023) Spatio-temporal Conditional Forecasting General BJTU SIGSPATIAL 2023
PriSTI (Liu et al., 2023c) Spatio-temporal Conditional Imputation General Beihang ICDE 2023
DiffUFlow (Zheng et al., 2023) Spatio-temporal Conditional Forecasting Transportation CSU CIKM 2023
DiffTraj (Zhu et al., 2024a) Spatio-temporal Conditional Generation Transportation SUSTech NeurIPS 2023
ControlTraj (Zhu et al., 2024b) Spatio-temporal Conditional Generation Transportation HKUST(GZ) arXiv 2024
Diff-RNTraj (Wei et al., 2024) Spatio-temporal Conditional Generation Transportation BJTU ArXiv 2024
Evans et al. (Evans et al., 2024) Univariate LDM Generation Audio Stability AI ArXiv 2024
LDT (Feng et al., 2024) Multivariate LDM Forecasting General NTU AAAI 2024
Aristimunha et al. (Aristimunha et al., 2023) Multivariate LDM Generation Healthcare UPSaclay NeurIPS 2023
TSDM (Pei et al., 2023) Multivariate DDIM Imputation&Anomaly Detection Electricity HUST ArXiv 2023
LADM (Lv et al., 2024) Spatio-temporal LDM Forecasting Transportation XJU IEEE T INSTRUM MEAS 2024
Diffusion Models for
Time Series and
Spatio-Temporal Data
Unconditional Diffusion Models (subsection 4.1 ) Probability-based Diffusion Models (Sec. 4.1.1) Predictive Tasks (subsection 5.1 subsection 5.4 ) Time Series (subsection 6.1 ) Healthcare: DiffEEG (Shu et al., 2023) DiffECG (Neifar et al., 2023) General: TimeGrad (Rasul et al., 2021) D3VAE (Li et al., 2022a) Spatio-Temporal Data (subsection 6.2 ) Climate: SwinRDM (Chen et al., 2023a) SEEDS (Li et al., 2023a) Traffic: DVGNN (Liang et al., 2023) SpecSTG (Lin et al., 2024) Generative Tasks (subsection 5.2 subsection 5.3 ) Time Series (subsection 6.1 ) Healthcare: Kuo et al. (Nicholas et al., 2023) Diff-E (Kim et al., 2023a) Audio: WaveGrad (Chen et al., 2020) DiffWave (Kong et al., 2020) Recommendation: DiffuASR (Liu et al., 2023g) DiffuRec (Li et al., 2023d) Spatio-Temporal Data (subsection 6.2 ) Traffic: TrajGDM (Chu et al., 2024) General: Yun et al. (Yun et al., 2023) Wang et al.(Wang et al., 2024b) Other Tasks Time Series (subsection 6.1 ) Classification: CARD (Han et al., 2022) Anomaly Detection: DDMT (Yang et al., 2023b) Spatio-Temporal Data (subsection 6.2 ) Anomaly Detection: DiffTAD (Li et al., 2024b) Imputation: Imputation as Inpainting (Yun et al., 2023) Score-based Diffusion Models (Sec. 4.1.2) Predictive Tasks (subsection 5.1 subsection 5.4 ) Time Series (subsection 6.1 ) General: ScoreGrad (Yan et al., 2021) Biloš et al. (Biloš et al., 2022) Spatio-Temporal Data (subsection 6.2 ) Climate: Hatanaka et al. (Hatanaka et al., 2023) General: DYffusion (Cachay et al., 2023) Generative Tasks (subsection 5.2 subsection 5.3 ) Time Series (subsection 6.1 ) Audio: SGMSE (Richter et al., 2023) StoRM (Lemercier et al., 2023) Healthcare: DeScoD-ECG (Li et al., 2023b) Spatio-Temporal Data (subsection 6.2 ) Healthcare: EHRDiff (Yuan et al., 2023c) Conditional Diffusion Models (Sec. 4.2.1) Probability-based Diffusion Models (Sec. 4.2.1) Predictive Tasks (subsection 5.1 subsection 5.4 ) Time Series (subsection 6.1 ) AIOps: Maat (Lee et al., 2023) Manufacture: Zhang et al. (Zhang et al., 2023e) SCDDPM (Chen et al., 2024a) Energy: VGCDM (Liu et al., 2023a) Dong at al. (Dong et al., 2023) Spatio-Temporal Data (subsection 6.2 ) Traffic: DiffUFlow (Zheng et al., 2023) DiffSTG (Wen et al., 2023) Climate: DiffESM (Bassetti et al., 2023) DiffSTG (Wen et al., 2023) Generative Tasks (subsection 5.2 subsection 5.3 ) Time Series (subsection 6.1 ) Healthcare: BioDiffusion (Li et al., 2024d) MEDiC (Sharma et al., 2023) Finance: DiffsFormer (Gao et al., 2024) Audio: CDiffuSE (Lu et al., 2022a) DOSE (Tai et al., 2024) Spatio-Temporal Data (subsection 6.2 ) Traffic: DiffTraj (Zhu et al., 2024a) ControlTraj (Zhu et al., 2024b) Others: PriSTI (Liu et al., 2023c) Score-based Diffusion Models (Sec. 4.2.1) Generative Tasks (subsection 5.2 subsection 5.3 ) Time Series (subsection 6.1 ) Healthcare: Seki et al. (Seki et al., 2023) Audio: UNIVERSE (Serrà et al., 2022)
Figure 6. A comprehensive taxonomy of diffusion models for time series and spatio-temporal data, categorized according to methodologies (i.e., unconditional vs. conditional), tasks (e.g., predictive versus generative), data types, and applications.
This section presents an overview and classification of diffusion models for addressing challenges in time series and spatio-temporal data analysis. Our survey organizes the discussion along four primary dimensions: categories of diffusion models, types of tasks, data modalities, and practical applications [^2] All summarized and categorized papers document: https://docs.google.com/spreadsheets/d/1X5ujA-yXCzkzz4Uu9ErP8mu_Jg2zWk4fwZrwrMcqMoI.. A comprehensive summary of notable related works is depicted in Figure 6 . We categorize the existing literature into two primary groups: unconditioned and conditioned diffusion models, focusing on time series and spatio-temporal data.
In the unconditioned category, diffusion models operate in an unsupervised manner to generate data samples without the need for supervision signals. This setting represents the foundational approach for analyzing time series and spatio-temporal data. Within this category, literature can be further divided into probability-based and score-based diffusion models. Examples include denoised diffusion probabilistic models (DDPMs) (Ho et al., 2020) and score-based stochastic differential equations (Score SDEs) (Song et al., 2020b; Song and Ermon, 2019), as introduced in section 2 . Research in this category is broadly organized into two task groups: predictive and generative tasks. Predictive tasks typically involve forecasting and anomaly detection, leveraging historical data and patterns to anticipate current and/or future events. Generative tasks, conversely, focus on identifying patterns within extensive datasets to generate new content, such as time series imputation and augmentation. Methods are developed for both primary data modalities: time series and spatio-temporal data, catering to a wide range of applications across various sectors, including healthcare, energy, climate, traffic, and more.
In the conditioned category, diffusion models are tailored for conditioned analysis of time series and spatio-temporal data. Empirical studies have shown that conditional generative models, which utilize data labels, are easier to train and yield superior performance compared to their unconditional counterparts (Bao et al., 2022). In this context, labels (a.k.a. conditions) often derive from various sources, such as extracted short-term trends (Shen and Kwok, 2023) and urban flow maps (Zheng et al., 2023), to enhance model inferences. This category embraces both probability-based and score-based diffusion models for predictive and generative tasks, offering a fresher perspective on leveraging diffusion models to tackle practical challenges in time series and spatio-temporal data analysis under specific constraints.
Building on the foundational understanding of model categories, task types, data modalities, and application domains, we delve deeper into the exploration of diffusion models for time series and spatio-temporal data analysis across the following sections. Each section is designed to unpack the complexities and nuances inherent in the application of diffusion models, providing a comprehensive overview from multiple perspectives. In section 4 , we explore diffusion model landscapes, highlighting distinctions between unconditioned and conditioned approaches and their implications. section 5 analyzes tasks from the predictive and generative viewpoints, detailing specific functions such as forecasting, generation, anomaly detection, and data imputation. section 6 examines the data modalities, differentiating between time series and spatio-temporal data to outline model challenges and applicability. Lastly, section 7 extends the discussion to application fields, demonstrating diffusion models’ utility across sectors like healthcare, traffic, sequential recommendation, climate, energy, and audio. This structured exploration aims to equip readers with an in-depth understanding of the potential and current state of diffusion models for addressing complex time series and spatio-temporal data challenges.
4. Model Perspective
In this section, we will analyze how to use diffusion models for time series and spatio-temporal data from a model perspective. Specifically, we will focus on standard diffusion models (DDPM and score SDE) and improved diffusion models (conditional diffusion model, LDM, DDIM, and others). In addition to the above content, we will also discuss in detail an important issue in the model section, namely, the denoisers in the diffusion model.
4.1. Standard Diffusion Model
Standard diffusion models include probability-based model, i.e., DDPM and score-based models, score SDE, which have been described in detail in Sec. 2.3.1 and Sec. 2.3.2, respectively. Currently, diffusion models based on these two are also the most common methods for time series and spatio-temporal data analysis.
4.1.1. Probability-Based Model
Based on the probabilistic-based standard model, we mainly introduce DDPM there, which uses a discrete framework, decomposing the diffusion process into a fixed number of steps. This method adds noise to the data at discrete intervals and then learns to reverse this process, generating data from noise. Both the training and inference process of the model are conducted in these discrete steps, making the process essentially stepwise and iterative.
TimeGrad is among the most classic DDPM-based time series forecasting methods, identified as a score-matching model that, through rigorous validation, has been proven effective in real-world datasets comprising thousands of interconnected dimensions (Rasul et al., 2021). D3VAE proposes a bidirectional variational auto-encoder that integrates diffusion, denoising, and disentanglement to enhance time series data without introducing extraneous uncertainties. This model marries multi-scale denoising score matching with bi-directional variational autoencoding for forecasting tasks (Li et al., 2022a). In a similar vein, TSDiff leverages implicit probability densities to iteratively refine the base forecasts, thereby addressing three distinct tasks: forecasting, refining, and generating synthetic data (Kollovieh et al., 2024). For the universal challenge of anomaly detection, Pintilie et al. have proposed two multivariate time series anomaly detection algorithms based on diffusion models, securing leading outcomes (Pintilie et al., 2023). Concurrently, D3R tackles temporal anomaly detection drifts through decomposition and reconstruction, employing data-time mix-attention for dynamic decomposition alongside diffusion models for end-to-end training, thus overcoming the limitations imposed by local sliding windows and unstable data scenarios (Wang et al., 2024c). CARD presents a denoising diffusion-based generative model alongside a pretrained conditional mean estimator to serve both classification and regression tasks (Han et al., 2022). TS-Diffusion is specifically designed for complex sequences marked by irregular sampling, missingness, and extensive temporal feature dimensions, introducing a comprehensive model consisting of an ODE encoder, a representational learning module, and an ODE decoder to handle such intricate time series (Li, 2023). Furthermore, WaveGrad (Chen et al., 2020), DiffWave (Kong et al., 2020), and DiffuASR (Liu et al., 2023g) have each successfully applied DDPM to waveforms, audio generation, and sequence recommendation, respectively.
On the other hand, in the field of spatio-temporal data, significant advances have been made with DDPM predominantly starting from 2023, indicating an exciting opportunity for more applications and theoretical breakthroughs. Yun et al. have put forward ”Imputation as Inpainting,” which integrates an unconditional diffusion model based on graph neural networks to first forecast complete spatio-temporal data. This approach adjusts the generation process by sampling in unobserved areas using the information from observed data, thus estimating missing values in spatio-temporal data (Yun et al., 2023). For traffic and specifically trajectory forecasting, the adoption of graph-based DDPM methods is becoming more common. For example, MID++ leverages a GNN-based ALEncoder to extract spatio-temporal features and improves the training process with an important sampling strategy, leading to better trajectory forecasting (Yao et al., 2023b). In a similar vein, DiffTAD incorporates a Transformer-based decoupled time and space encoder to model spatial interactions among vehicles and conducts anomaly detection by evaluating the differences between query trajectories and their reconstructions (Li et al., 2024a). Moreover, SpecSTG tackles the challenges of insufficient representation of spatial network features and the inability to detect unexpected fluctuations in future observations in traffic flow forecasting. It achieves this by converting the learning process to the spectral domain, generating Fourier representations of future time series that carry spatial information (Lin et al., 2024).
4.1.2. Score-Based Model
Unlike the discrete DDPM, score-based SDE models represent diffusion and the reverse process in a continuous form using stochastic differential equations, thereby covering continuous time. This approach allows for a theoretically more flexible and in-depth treatment of the diffusion process, capable of generating samples at any point in time, rather than being confined to fixed steps.
ScoreGrad and TimeGrad share similar goals and can be described as SDE-based TimeGrad models. They expand the diffusion process into a continuous spectrum and employ time series feature extraction modules along with a condition-based stochastic differential equation for score matching to facilitate forecasting (Yan et al., 2021). This structure finds parallels in the work of Biloš et al. (Biloš et al., 2022). Additionally, recent studies by Crabbé et al. delve into representing time series data in the frequency domain using SDEs, showcasing how denoising score matching methods can facilitate diffusion modeling within the frequency domain based on differing time-frequency spaces (Crabbé et al., 2024). TimeADDM introduces a score-based diffusion model for unsupervised anomaly detection in multivariate time series, applying diffusion steps to representations that encapsulate global time correlations through recurrent embeddings and designing a suite of reconstruction strategies to compute anomaly scores at various diffusion intensities (Hu et al., 2024). Moreover, Li et al. extensively discuss the impact of noise samples in SDEs on data quality and the corresponding robustness of the models (Li et al., 2024c).
From an application standpoint, score-based diffusion models have shown great promise in continuous data fields such as healthcare and acoustic data. EHRDiff (Yuan et al., 2023c) and DeScoD-ECG (Li et al., 2023b) leverage score-based diffusion models for realistic EHR synthesis and ECG data generation, respectively, incorporating algorithms or structures pertinent to the features of medical data, thus improving applicability in real-world settings. In audio-related applications, the focus has been on speech enhancement and dereverberation, with initial proposals by Welker et al. (Welker et al., 2022) and Richter et al. (Richter et al., 2023) for employing score-based diffusion models for speech enhancement, transforming time domain speech data into time-frequency information, and treating them as image input for models. StoRM introduces a stochastic resampling method, utilizing forecasting from a predictive model as guidance for subsequent diffusion steps, achieving higher-quality samples under complex conditions with high signal-to-noise ratios using fewer diffusion steps and more streamlined models (Lemercier et al., 2023). Similarly, Lay et al. emphasize improving model efficiency and optimizing model parameters, proposing a Brownian bridge-based forward process to reconcile the gap between the forward process’s end distribution and the prior distribution used in inference for the reverse process (Lay et al., 2023).
In the realm of spatio-temporal data, methods remain scarce, with Sasdim and DYffusion standing out. Sasdim, an adaptive noise scaling diffusion model, effectively performs spatio-temporal data imputation by capturing dynamic spatio-temporal dependencies through novel loss functions and global spatio-temporal convolution modules (Zhang et al., 2023d). DYffusion, a dynamics-informed diffusion model for spatio-temporal data forecasting, distinguishes itself from probabilistic models by integrating temporal dynamics directly into the diffusion steps and training a stochastic, time-conditioned interpolator alongside a predictor network to simulate the forward and reverse processes of the standard diffusion model, thus striking a balance between performance and efficiency (Rühling Cachay et al., 2024).
4.2. Improved/Advanced Diffusion Model
We have already discussed the improvement methods of the diffusion model in Sec. 2.3.3 and Sec. 2.3.4. In the following, we will analyze the methods of the more frequently used ones.
4.2.1. Conditional Diffusion Model
Compared to standard approaches, conditional diffusion models are able to use given conditional information (e.g., different representations, different modalities, etc.) to steer the generative process and produce high-quality outputs that are tightly correlated with conditions (Shen and Kwok, 2023). More condition-based models are used for high-quality generation tasks.
CSDI and MIDM use conditional diffusion models for time series imputation (Tashiro et al., 2021; Wang et al., 2023d). They utilize score-based and probability-based diffusion models conditioned on observed values and through explicit imputation training. They can leverage the correlations among observed values to further enhance the performance. Similar to CSDI, DiffAD and ImDiffusion employ a similar method based on conditional weight-incremental diffusion to enhance the imputation performance of missing values and are applied for time series anomaly detection (Xiao et al., 2023; Chen et al., 2023c). These approaches preserve the information on the observed values and significantly improve the generation quality for stable anomaly detection. On the other hand, there are models that utilize different guiding information (Fu et al., 2024). For example, (Zuo et al., 2023) adjusts the diffusion model based on statistical information such as mean, standard deviation, Z-scores, and skewness, thereby synthesizing sensor data. (Tai et al., 2024) proposes two different conditional enhancement techniques to enable the model to adaptively consider conditional information a priori, thus performing the speech enhancement task. MEDiC (Sharma et al., 2023) introduces a class-conditional DDPM approach to generate synthetic EEG embeddings. VGCDM (Liu et al., 2023a) employs a pulse voltage-guided diffusion model along with a cross-attention mechanism for more efficient generation of electrical signals. Meanwhile, Wang et al. leverage multimodal information (images and text) as generative conditions to enhance the quality of real-time sales forecasts (Wang, 2024). DiffShape (Liu et al., 2024b) introduces a self-supervised diffusion learning mechanism that uses real sub-sequences as conditions. By leveraging a large amount of unlabeled data, it enhances the similarity between learned shapelets and actual sub-sequences for classification.
For spatio-temporal data, most models adopt graph structures for representation, but the method of using conditions is similar to that of time series data (Ling et al., 2024; Liu et al., 2024a). PriSTI draws inspiration from CSDI of conditioning on the observed value, employing a conditional feature extraction module based on linear interpolation for generating missing data (Liu et al., 2023c). DiffTraj estimates noise levels accurately using various external factors, such as the region of the trip and departure time (Zhu et al., 2024a). ControlTraj further extends DiffTraj with the constraint of road network structures (Zhu et al., 2024b). Additionally, many works guide their models using graph structures or feature graphs as conditions. For instance, DiffSTG (Wen et al., 2023) and USTD (Hu et al., 2023) use historical graph signals and graph structures as conditions, combining the spatio-temporal learning capabilities of GNNs and the uncertainty measurement of diffusion models to enhance the performance of forecasting tasks. DiffUFlow (Zheng et al., 2023) utilizes extracted spatio-temporal feature graphs overlaid on coarse-grained flow graphs as a condition to guide the reverse process.
4.2.2. Latent Diffusion Model
The latent diffusion model (LDM) performs the diffusion process in a lower-dimensional latent space, which allows more efficient generation or other tasks. It allows the model to handle more complex data distributions, while reducing computational resource consumption and maintaining output quality.
LDCast utilizes LDM for near-term precipitation forecasting, enhancing training stability, optimizing computational demands, and accurately representing the uncertainty of forecasting compared to GAN (Leinonen et al., 2023). CLDM presents a conditional latent diffusion model based on latent space mapping, which breaks down the generative task into deterministic forecasting and the generation of predictive error scenarios, thus increasing efficiency within the latent space (Dong et al., 2023). Similarly, Aristimunha et al. suggest the use of LDM for generating sleep EEG signals, where the latent feature maps outputted by the encoders serve as input to the diffusion model, with the generated results obtained through the corresponding decoder (Aristimunha et al., 2023). In a similar manner, LADM proposes a method for spatio-temporal trajectory forecasting, where VAE acts as a generator and DDPM as a refiner (Lv et al., 2024). Furthermore, Feng et al. develop a latent diffusion transformer for time series forecasting, aimed at compressing multivariate time stamp patterns into succinct latent representations and efficiently generating authentic multivariate time stamp values in a continuous latent space (Feng et al., 2024).
4.2.3. Other Variants of Diffusion Models
There are also other methods that incorporate the different models above or use other more advanced diffusion models. For example, to accelerate generation, MedDiff for the first time utilizes DDIM with a new sampling strategy to generate high-dimensional large-scale electronic medical records (He et al., 2023). Specifically, it incorporates a classifier to guide the sampling process, assigning high probabilities to the data with the correct labels. DCM is a diffusion-based causal inference model accelerated by DDIM, that more accurately captures counterfactual distributions under unmeasured confounding factors (Shimizu, 2023). Asperti et al. propose GED, which uses conditional DDIM integrated with additional information and a post-processing network for high-performance weather forecasting (Asperti et al., 2023).
Besides, some algorithms have also been proposed that use a two-stage or a diffusion model to fuse with other models. TSDM adopts an improved two-stage diffusion model for identifying and reconstructing measurements with various uncertainties for power system measurement recovery, where the first stage includes a classifier-guided conditional anomaly detection component, and the second stage involves a diffusion-based measurement imputation component (Pei et al., 2023). Mueller proposes using attention-enhanced condition-guided DDIM to tackle sample imbalance and data scarcity issues, along with simple DDPM for signal denoising, thereby synthesizing effective machine fault data for efficient anomaly detection (Mueller, 2024). Similarly, Wong et al. utilize DDIM for load domain adaptation, further enhanced with CNN for bearing fault anomaly detection (Wong et al., 2024).
4.3. Denoisers in the Diffusion Model
Denoisers in the diffusion model serve as the core modules that transform noise into structured signals, with their architectures, inductive biases, and prior-guided designs fundamentally shaping the model’s generative capacity, temporal reasoning, and applicability to diverse inverse problems. In the following, we will provide a detailed analysis of these aspects respectively.
4.3.1. Architecture of the Denoiser
The denoiser serves as the backbone of diffusion models, determining their expressive power and cross-modal generalization. U-Net remains the architecture most widely used in diffusion models (Ronneberger et al., 2015; Ho et al., 2020; Rombach et al., 2022), where its encoder–decoder design with skip connections supports multi-scale representation learning and effective propagation of global and local information. Models such as DiffuASR (Liu et al., 2023g) and DiffAD (Xiao et al., 2023) adapt U-Net with 1D convolutions or temporal dilations to better model sequential dependencies. Some works explore attention-based backbones for long-range dependency modeling, including CSDI (Tashiro et al., 2021) and PriSTI (Liu et al., 2023c), which enhance global context capture and scalability (Zheng et al., 2023). Hybrid architectures further extend denoising to structured data. D3R combines attention and diffusion for anomaly detection (Wang et al., 2024c), while SpecSTG uses graph spectral convolutions for traffic forecasting (Lin et al., 2024). These domain-aware designs show that integrating prior knowledge improves denoising efficiency and interpretability (Wen et al., 2023). Recent trends favor modular, cross-modal and lightweight denoisers, including foundation backbones with temporal or structural adaptations (Ma et al., 2025).
4.3.2. Inductive Bias in Temporal Denoisers
The effectiveness of diffusion models relies strongly on the inductive biases in their denoiser architectures. U-Net structures leverage translation equivariance and hierarchical feature extraction for robust spatial representations (Kadkhodaie et al., 2024; Kamb and Ganguli, 2025). For time series, analogous inductive priors must capture temporal continuity, periodicity, and causal dynamics. Hierarchical temporal architectures (e.g., Temporal U-Net, WaveGrad (Chen et al., 2020)) address multi-scale temporal patterns through progressive downsampling and upsampling. In spatio-temporal settings, CNN-based denoisers (e.g., DiffTraj (Zhu et al., 2024a)) learn long-term dependencies efficiently with skip connections, while graph-based denoisers (e.g., DiffSTG (Wen et al., 2023), DiffUFlow (Zheng et al., 2023)) incorporate structural priors via graph convolutions. Emerging temporal transformers (e.g., DiffUFlow (Zheng et al., 2023), PriSTI (Liu et al., 2023c)) unify causal and hierarchical biases through self-attention with relative positional encodings.
4.3.3. Denoiser-as-Prior in Temporal Inverse Problems
Beyond the full diffusion sampling framework, many inverse problems can be solved by directly leveraging the denoiser (or score function) as a learned prior. This idea has been widely explored under the Plug-and-Play and Score Prior paradigms, where the denoiser acts as a proximal operator for iterative reconstruction (Kadkhodaie and Simoncelli, 2020; Milanfar and Delbracio, 2025). For the time series field, a similar trend has emerged. Models such as ImDiffusion (Chen et al., 2023c) and SADI (Islam et al., 2025) employ the denoiser alone to restore missing segments, achieving high-quality reconstructions without full stochastic sampling. This approach reduces computational overhead and enhances interpretability by decoupling temporal prior learning from generative sampling. The denoiser-as-prior framework therefore provides a promising bridge between generative diffusion modeling and classical temporal signal restoration.
5. Task Perspective
In this section, we will explore the use of diffusion models in different tasks, such as forecasting, generation, imputation, and anomaly detection, also highlighting their effectiveness in complex time series and spatio-temporal data analysis across various domains.
5.1. Forecasting
The field of time series forecasting has seen significant advancements with the incorporation of diffusion models. TimeGrad (Rasul et al., 2021) and D3VAE (Li et al., 2022a) both employ diffusion probabilistic models to enhance forecasting, with TimeGrad focusing on autoregressive techniques (i.e., RNN) for probabilistic forecasting and D3VAE introducing a bidirectional variational auto-encoder that includes diffusion and denoising processes. This generative approach is further extended in the work by D-Va (Koa et al., 2023), which addresses the stochastic nature of stock price data through a deep hierarchical VAE combined with diffusion probabilistic techniques. Meanwhile, the study presented in (Biloš et al., 2022) takes a different approach by modeling temporal data as continuous functions, allowing the handling of irregularly sampled data. Building upon the diffusion model, TimeDiff (Shen and Kwok, 2023) introduces novel conditioning mechanisms, future mixup, and autoregressive initialization, to improve time series forecasting. Lastly, the research in (Kollovieh et al., 2024) explores task-agnostic unconditional diffusion models, proposing TSDiff, which employs a self-guidance mechanism for versatile time series applications.
For spatio-temporal data, various diffusion model-based approaches have been proposed to tackle complex forecasting problems. DiffSTG (Wen et al., 2023) is the first work that generalizes denoising diffusion probabilistic models to spatio-temporal graphs, aiming to model complex spatio-temporal dependencies and intrinsic uncertainties within spatio-temporal graph data for better forecasting. DYffusion (Rühling Cachay et al., 2024) introduces a framework that trains a stochastic, time-conditioned interpolator and a forecaster network to perform multi-step and long-range probabilistic forecasting for spatio-temporal data. On the other hand, DiffUFlow (Zheng et al., 2023) focuses on urban data, which aims to address the challenge of fine-grained flow inference. DSTPP (Yuan et al., 2023a) provides a novel parameterization for spatio-temporal point processes. Meanwhile, SwinRDM (Chen et al., 2023a) and DOT (Lin et al., 2023b) demonstrate the adaptability of diffusion models in improving the quality of weather forecasts and travel time estimations, showcasing the wide applicability of these models across different domains and tasks. These works highlight the growing impact and potential of diffusion models in advancing spatio-temporal data analysis.
5.2. Generation
Inspired by the powerful ability of the diffusion model in high-dimensional distribution learning, an intuitive usage is applying the learned diffusion model for data generation. Currently, a variety of diffusion models have been proposed for generating audio data. WaveGrad (Chen et al., 2020) is a pioneering work that utilizes gradient-based sampling to generate high-fidelity audio waveforms, marking a significant advancement in audio synthesis. Similarly, DiffWave (Kong et al., 2020) employs a non-autoregressive approach to achieve efficient and high-quality raw audio synthesis through a Markov chain process. Both WaveGrad and DiffWave are part of the same research lineage, leveraging the power of diffusion models to create complex waveforms from simple noise distributions. DOSE (Tai et al., 2024) takes a different approach by focusing on the speech enhancement task, which introduces a model-agnostic method that integrates condition information into the diffusion process. Besides, diffusion model is also used in sequential recommendation. DiffuASR (Liu et al., 2023g) proposes a diffusion-based sequence generation framework to address data sparsity and long-tail user problems in sequential recommendation systems. It introduces a sequential U-Net designed for discrete sequence generation tasks and utilizes two guide strategies to assimilate preferences between generated and original sequences. DreamRec (Yang et al., 2024b) employs a Transformer encoder to create guidance representations as conditions in the diffusion process.
In the realm of spatio-temporal data, diffusion models are adopted to generate trajectory data. For instance, DiffTraj (Zhu et al., 2024a, b) proposes the first effort that generates high-quality human trajectories with an unconditioned model, with the motivation to protect privacy. Meanwhile, (Lin et al., 2023b) generates trajectories in the format of grids, by a condition model guided by the origin-destination information.
5.3. Imputation
In the domain of time series and spatio-temporal data analysis, imputation refers to generating the unobserved data conditioned on the given observed data. CSDI (Tashiro et al., 2021) proposes a score-based diffusion model that probabilistically imputes missing time series and spatio-temporal data. MIDM (Wang et al., 2023d) redefines the evidence lower bound (ELBO) for conditional diffusion models tailored for multivariate time series imputation, which ensures the consistency of observed and missing values. PriSTI (Liu et al., 2023c) introduces a conditional diffusion framework specifically designed for spatio-temporal data imputation, using global context priors and geographic relationships to tackle scenarios with high missing data rates due to sensor failures. More recently, TabCSDI (Zheng and Charoenphakdee, 2022) and MissDiff (Ouyang et al., 2023) leverage diffusion models for imputing missing values in tabular data, each addressing distinct aspects of this complex challenge. TabCSDI manages categorical and numerical data effectively, providing a tailored solution to diverse data types common in tabular datasets. On the other hand, MissDiff enhances the diffusion model’s training process by introducing a novel masking technique for regression loss, which ensures consistent learning of data distributions and robustness against different missing data scenarios.
5.4. Anomaly Detection
Anomaly detection aims to identify anomalies in the given time series or spatio-temporal data, which is quite practical and crucial for many real-world applications. A majority of works have been proposed for general time series anomaly detection. Initially, DiffAD (Xiao et al., 2023) and ImDiffusion (Chen et al., 2023c) both explored the synergy of imputation techniques with diffusion models for time series anomaly detection, enhancing the robustness of anomaly detection processes by accurately modeling complex dependencies. Concurrently, (Pintilie et al., 2023) and (Livernoche et al., 2023) use similar diffusion-based methodologies to address anomaly detection but apply different enhancements to improve performance and computational efficiency. Meanwhile, (Wang et al., 2024c) and DDMT (Yang et al., 2023b) tackled the specific challenges of instability and noise, implementing advanced diffusion reconstruction techniques to maintain accuracy in dynamic environments.
Besides the above methods designed for general time series anomaly detection. There are other works that apply advanced diffusion techniques across various domains. For instance, Maat (Lee et al., 2023) anticipates performance metric anomalies in cloud services using a conditional denoising diffusion model to forecast metrics and detect anomalies. (Mueller, 2024) enhances data synthesis for machine fault diagnosis using an attention mechanism within a conditional diffusion model framework. Diffusion-UDA (Zhao et al., 2024c) proposes a diffusion-based method for unsupervised domain adaptation in submersible fault diagnosis, leveraging diffusion processes to adapt domains for effective fault recognition.
6. Data Perspective
In this section, we analyze current work from a data perspective, including time series and spatio-temporal data, as shown in Figure 4 . We focus on introducing how diffusion models capture the unique properties of different data modalities.
6.1. Time Series
Time series analysis is a critical issue explored across various real-world scenarios, including retail sales forecasting, filling in missing data in economic time series, identifying anomalies in industrial maintenance, and categorizing time series from different domains.
6.1.1. Univariate Time Series
Univariate time series is characterized by having only one variable of interest observed over a period of time. Data in this category includes ECG signal, audio, electricity load, etc. The complex sequential patterns (e.g., the trend and the periodicity) are the core data property in univariate time series, which pose great challenges in developing relevant diffusion models.
Diffusion models for univariate time series are primarily developed to model the uncertainty present in the data, facilitating tasks such as probabilistic forecasting or data generation. An essential element within the diffusion model is the denoising network, which determines the extent of noise removal at each stage. In the context of univariate time series, the denoising network commonly employs the 1-D CNN to identify sequential patterns in the input data. For instance, both WaveGrad (Chen et al., 2020) and DiffWave (Kong et al., 2020) incorporate a 1-D CNN into the denoising network to extract local sequential features from audio sequences. Furthermore, some studies opt to transform the univariate time series into the frequency domain initially to enhance the capture of long-term sequential correlations from a global standpoint. This transformation converts the time series from 1-D to 2-D data, enabling the use of a 2-D CNN in the denoising network to capture correlations within the frequency domain, such as (Koa et al., 2023) for the stock price forecasting and DiffLoad (Wang et al., 2023b) for electricity load forecasting.
6.1.2. Multivariate Time Series
Multivariate time series is characterized by having multiple variables of interest observed over the same period of time. Instead of the sequential patterns in each variate, multivariate time series also leverage the interdependencies among different variables to capture more comprehensive information for downstream tasks.
Beyond approaches that focus on univariate time series, there have been efforts towards multivariate time series. Multivariate time series is naturally a 2-D data similar to image data, where the signal of each variate can be considered as one channel in the image. To this end, diffusion models for multivariate time series usually adopt the vanilla U-Net as the denoising net, which is a common practice in diffusion models for images.
For example, (Durall et al., 2023) first transforms demultiplexing, denoising, and interpolation into image-to-image transformation tasks, then leverages the diffusion model for different tasks. In the demultiplex task, the diffusion model learns to remove the multiples without removing the primary energy. In the denoising and interpolation task, the diffusion model learns to eliminate undesired uncorrelated noise and recover the missing values, while preserving the inherent characteristics of the data. MIDM (Wang et al., 2023d) proposes a novel multivariate imputation diffusion model that incorporates correlations between observed and missing values. By re-deriving the ELBO of the conditional diffusion model, MIDM ensures consistency between observed and missing data, thus leading to improved imputation accuracy. Diff-E (Kim et al., 2023b) further utilizes the diffusion model for representation learning, which combines DDPMs with a conditional autoencoder to enhance the decoding performance of speech-related EEG signals.
6.2. Spatio-Temporal Data
In real-world systems, a myriad of elements interact with each other both spatially and temporally, resulting in a spatio-temporal composition. Spatio-temporal data (STD) is the de facto most popular data structure for injecting such structural information into the formulation of practical problems. In this section, we introduce developments of diffusion models for spatio-temporal data, mainly with two data modalities: (1) the Spatio-Temporal Graph (STG) and (2) the Spatio-Temporal Trajectory (STT). Unlike diffusion models for time series data, existing works in STD need to model the temporal dependencies as well as the spatial dependencies, making the task even more challenging.
6.2.1. Spatio-Temporal Graph
The spatio-temporal graph is generated by sensors at different places in a period of time, where the correlation of those sensors is typically described as a graph. We categorize diffusion models for spatio-temporal graphs into domain-oriented and domain-agnostic works.
Most of the existing works fall into the category of domain-oriented spatio-temporal graph diffusion models. They typically leverage the powerful distribution learning abilities of diffusion models for spatio-temporal data mining tasks in specific domains, such as traffic and climate. In this research trend, the majority of works come from the traffic domain. DiffUFlow (Zheng et al., 2023) represents pioneering efforts as traffic diffusion models, which convert fine-grained urban flow inference into a denoising diffusion process. SpecSTG (Lin et al., 2024) advances the field by conducting the diffusion process in the spectral space projected by the STG, resulting in faster inference speeds. In a similar vein, (Liang et al., 2023) uses a diffusion model to infer the link probability and reconstruct causal graphs in the decoder stage adaptively for the STG forecasting task. Furthermore, (Lin et al., 2023b) solves the origin-destination travel time estimation problem with a two-step process, where the first step predicts the possible travel route with a diffusion model conditioned on the given origin-destination pair. Beyond the traffic domain, there have been initiatives towards the climate. A notable example is (Chen et al., 2024b), which quantifies the uncertainty of air quality forecasting based on the spatio-temporal diffusion model. Similarly, SRNDiff (Ling et al., 2024) conducts the short-term rainfall nowcasting through the conditional diffusion model.
In the second group, researchers focus on developing domain-agnostic models, which can achieve promising performance across a variety of domains. For instance, DiffSTG represents a pioneering effort by introducing a unified diffusion framework for multiple STG tasks, such as forecasting and imputation. Concurrently, PriSTI (Liu et al., 2023c) and (Yun et al., 2023) study the spatio-temporal imputation task while treating the task of recovering unobserved values as a denoising process conditioned on the observed values. Similarly, DYffusion (Cachay et al., 2023) models the temporal dynamics directly within diffusion steps, leading to a stochastic, time-conditioned forecasting network. Another noteworthy contribution is (Hu et al., 2023) under the domain-agnostic and task-agnostic setting, which aims to harness the power of diffusion models, and unify diffusion models for probabilistic spatio-temporal graph learning.
6.2.2. Spatio-Temporal Trajectory
Spatio-temporal trajectory is a sequence of locations ordered by time that describes the movements of an object in a geographical space. Analysis of spatio-temporal modality is particularly crucial to discovering the mobility patterns of moving objects, which serves as the foundation for many downstream tasks, such as POI recommendation and next-location forecasting. In terms of diffusion models for trajectory data, most works are developed to solve the trajectory generation task. They leverage the ability of the diffusion model to learn high-dimensional data distribution while injecting spatio-temporal correlation into the diffusion process. As exemplified by DiffTraj (Zhu et al., 2024a), which generates the GPS trajectory with the diffusion probabilistic model in an unconditioned manner. Similarly, (Chu et al., 2024) utilizes the diffusion model for generating high-quality human mobility data. More recently, Diff-RNTraj (Wei et al., 2024) further generates the trajectory constrained by the road-network. Beyond generation, there is a growing interest in applying diffusion models for trajectory forecasting tasks, and (Yao et al., 2023a) and (Liu et al., 2024a) exemplify this trend.
7. Application Perspective
In this section, we summarize and discuss the most important applications of diffusion models in time series and spatio-temporal data, including healthcare, smart city, recommendation, climate and weather, energy and electricity, audio, video, and so on. The toy signal examples of each application are shown in  Figure 7 .
Figure 7. Examples of time series and spatio-temporal data for different application scenarios.
In detail, we will discuss the models in different applications, aiming to show light on the model design of practical applications.
7.1. Healthcare
In recent years, diffusion models have emerged as a powerful class of generative models with significant implications for time series in healthcare. These models, known for their ability to generate high-fidelity samples through a process of gradual refinement, have found diverse applications ranging from the synthesis of electronic health records (EHRs) to the enhancement of biomedical signal analysis. The following are some main topics in this area. EEG Signal Synthesis and Enhancement: (Neifar et al., 2023) introduces a generalized probabilistic diffusion model tailored for the synthesis of electrocardiogram (ECG) signals, aiming to support cardiac health research and diagnostic training with realistic data generation. (Aristimunha et al., 2023) demonstrates the potential of diffusion models to create synthetic EEG signals, facilitating sleep disorder studies and neurological research. (Li et al., 2023b; Alcaraz and Strodthoff, 2023) focus on enhancing the quality of ECG signals, addressing issues like baseline wander and noise, and generating conditional ECG signals based on specific patient states. Healthcare Data Augmentation and Synthesis: (Shu et al., 2023) highlights the use of diffusion models to augment datasets for seizure forecasting, thereby improving the robustness and accuracy of predictive models. (Yuan et al., 2023c; Ceritli et al., 2023) explore the synthesis of realistic EHRs, tackling the challenges of data privacy and scarcity by providing synthetic datasets for research and model training. Diagnostic and Predictive Analytics: (Li et al., 2023c) leverages generative models to enhance the detection of autism, showcasing the broad applicability of these models in diagnosing and understanding complex conditions. In (Chang et al., 2024), diffusion models are applied to forecast critical physiological parameters in ICUs, demonstrating their potential to save lives by predicting adverse events before they occur. Novel Methodologies and Techniques: (Zhao et al., 2024b) exemplifies the versatility of diffusion models in handling multiple tasks simultaneously, such as predicting mortality in critically ill patients, showcasing the potential of comprehensive care management systems.
7.2. Traffic
There are also many works for traffic applications using diffusion models. The following are some of them. DiffUFlow (Zheng et al., 2023) employs a denoising diffusion model to enable the model to capture the stochastic nature of urban flows and generate accurate and fine-grained forecasting. DiffTraj (Zhu et al., 2024a) presents a groundbreaking approach to generating GPS trajectories using a spatio-temporal diffusion probabilistic model. The model works by reconstructing geographic trajectories from white noise through a reverse trajectory denoising process, effectively turning random noise into meaningful trajectory data that reflects real-world movement patterns. DVGNN architecture (Liang et al., 2023) comprises an encoder component that learns latent node embeddings through GCN layers and a decoder component that treats the relationships of latent states as a diffusion process subject to stochastic differential equations. This allows for the inference of internal causal relationships among neighbor nodes and the formation of dynamic causal graphs. (Guimarães et al., 2023) addresses the growing risk of collisions between space objects by developing a diffusion model that predicts the positional uncertainty of objects during close encounters. (Li et al., 2024a) introduces a model that utilizes denoising diffusion probabilistic models for identifying anomalies in vehicle trajectories, aimed at improving the detection of unusual patterns in vehicle movements. SpecSTG (Lin et al., 2024) aims to efficiently handle the complexities of spatio-temporal data, offering a probabilistic approach to forecasting traffic flows. This method is distinguished by its speed and accuracy in generating forecasts and is a promising solution for real-time traffic management and planning applications.
7.3. Sequential Recommendation
In the scenario of sequential recommendation, diffusion models are also widely applied. (Liu et al., 2023g) and (Wang et al., 2023c) both explore the application of diffusion models in enhancing sequential recommendation systems, yet they focus on different aspects and methodologies within this domain. That is, (Liu et al., 2023g) aims to bridge the gap between discrete item identities and the continuous nature of the data generated by the diffusion models, while (Wang et al., 2023c) focuses on predicting users’ future interaction probabilities by corrupting and then denoising their interaction histories and addressing challenges specific, such as high resource costs and temporal shifts in user preferences. DCDR (Lin et al., 2023a) is a discrete forward process with tractable posteriors and a conditional reverse process tailored for sequence generation. RecFusion (Bénédict et al., 2023) introduces a binomial diffusion process tailored for one-dimensional data, such as time series of user interactions, showcasing its effectiveness in sequential recommendation scenarios. DiffuRec (Li et al., 2023d) proposes a specific diffusion model framework designed for sequential recommendation, focusing on efficiently capturing and predicting evolving user interests. Diffusion models have shown significant promise in improving sequential recommendation systems. Their ability to simulate the spread of information and preferences offers a nuanced method for predicting user behavior. Future research could explore hybrid models, scalability, and real-time adaptation, further enhancing the relevance and personalization of recommendations.
7.4. Climate and Weather
Diffusion models, initially developed for image generation, have found novel applications in weather forecasting due to their ability to generate high-resolution, realistic outputs. These models work by gradually refining a signal from a random noise distribution towards a desired output, making them well-suited for predicting complex atmospheric phenomena. (Leinonen et al., 2023) focuses on precipitation nowcasting with an emphasis on accurate uncertainty quantification, leveraging latent space representations to efficiently model and predict rainfall intensity and distribution. (Kurinchi-Vendhan, [n. d.]) presents a wind resolution-enhancing model that improves the detail and accuracy of wind speed forecasts by refining coarse forecasting to higher resolutions. SwinRDM (Chen et al., 2023a) integrates Swin Recurrent Neural Networks (SwinRNN) with diffusion models for state-of-the-art weather forecasting, achieving high resolution and quality by capturing both spatial and temporal dynamics. DiffESM (Bassetti et al., 2023) utilizes diffusion models for the conditional emulation of Earth System Models (ESMs), offering a computationally efficient alternative to traditional simulation techniques with improved accuracy. (Hatanaka et al., 2023) addresses the challenge of high-resolution solar forecasting, showcasing the model’s ability to accurately predict solar irradiance variations. DiffAD (Huang et al., 2024) introduces a diffusion model-based approach for weather-scale data assimilation, enhancing the integration of observational data into forecast models for improved accuracy.
7.5. Energy and Electricity
Diffusion models are now being adapted to tackle challenges in energy systems, offering novel solutions for probabilistic forecasting, signal synthesis, and system security. DiffLoad (Wang et al., 2023b) is a novel approach to load forecasting that leverages diffusion models to quantify uncertainties effectively. By incorporating uncertainty quantification, DiffLoad offers a more reliable forecast, which is crucial for grid stability and operational planning. (Wang and Zhang, 2023) explores how conditional diffusion models can synthesize load profiles for individual electricity customers, enhancing the personalization of demand-side management strategies. DiffPLF (Li et al., 2024f) addresses the challenges of load forecasting by providing probabilistic forecasts of EV charging loads, facilitating better grid management and infrastructure planning. (Capel and Dumas, 2023) discusses the application of denoising diffusion probabilistic models in energy forecasting, emphasizing the model’s ability to handle uncertainty and provide probabilistic forecasts, which are essential for integrating renewable energy sources. (Dong et al., 2023) demonstrates how conditional latent diffusion models can be used to generate realistic short-term wind power scenarios, aiding in system operation and planning.
7.6. Audio
Diffusion models have also found promising applications in generating waveforms, synthesizing audio, music generation and enhancing speech, offering significant improvements in quality, realism, and control over the generated or enhanced audio. Waveform Generation and Audio Synthesis: WaveGrad (Chen et al., 2020) introduces a gradient-based approach to generate high-fidelity waveforms, demonstrating the potential of diffusion models in audio synthesis without the need for autoregressive models. DiffWave (Kong et al., 2020) extends the capabilities of diffusion models in audio synthesis, showcasing their versatility across different audio synthesis tasks, including voice and music. Music Generation: (Zhang et al., 2023a) explores the fusion of diffusion models and GANs for generating symbolic music, allowing for emotion-driven control over the generation process. DiffuseRoll (Wang, 2023) highlights the application of diffusion models in multi-track music generation, providing nuanced control over various musical attributes. Speech Enhancement: (Richter et al., 2023) introduces a diffusion-based generative approach to simultaneously enhance speech and reduce reverberation, showcasing significant improvements over traditional methods. (Serrà et al., 2022) focuses on the universal applicability of diffusion models and demonstrates effectiveness in enhancing speech across a wide range of conditions. (Lay et al., 2023) proposes a method to reduce the prior mismatch in stochastic differential equations, enhancing model performance in speech tasks. CRA-DIFFUSE (Qiu et al., 2023) introduces a pre-denoising step in the time-frequency domain to improve cross-domain speech enhancement, illustrating the potential for methodological innovations within diffusion model applications. The application of diffusion models in audio processing has opened new avenues for research and development, offering novel solutions to longstanding challenges. As the field progresses, future research may focus on improving model efficiency, reducing computational demands, and exploring untapped applications within audio processing.
7.7. AIOps
Cloud computing’s reliability is paramount, yet it faces challenges like performance anomalies, unpredictable network traffic, and incomplete data. Diffusion models have emerged as a powerful tool to tackle these issues, offering new approaches for predictive maintenance, realistic traffic simulation, and enhanced failure forecasting. Maat (Lee et al., 2023) applies conditional diffusion models to anticipate performance metric anomalies in cloud services, demonstrating the potential for early anomaly detection, potentially reducing downtime, and improving service reliability. NetDiffus (Sivaroopan et al., 2023) applies diffusion models to time-series imaging for generating realistic network traffic patterns, showing how simulated traffic can support network planning, testing, and anomaly detection, enhancing network management and security. (Yang et al., 2023c) investigates diffusion models for imputing missing data in time series, aiming to improve cloud failure forecasting accuracy, highlighting the effectiveness of diffusion models in handling data gaps, leading to better predictive outcomes for cloud service failures.
7.8. Finance
The complexity and stochastic nature of financial markets make them an ideal candidate for the application of diffusion models, which can capture non-linearities and intricate patterns in data. Recent advancements have highlighted the potential of diffusion models in enhancing stock price forecasting, generating realistic financial tabular data, and augmenting stock factor data for improved investment strategies. (Koa et al., 2023) presents a novel approach to predict stock prices using a Diffusion Variational Autoencoder, aiming to better capture the stochastic nature of the market, highlighting the model’s ability to handle market volatility. FinDiff (Sattarov et al., 2023) focuses on generating synthetic financial tabular data to address the scarcity of publicly available financial datasets. The synthetic data generated by FinDiff closely mimics real financial datasets, potentially aiding in model training and regulatory compliance without compromising sensitive information. DiffSTOCK (Daiya et al., 2024) employs diffusion models for probabilistic relational reasoning in stock market forecasting, aiming to capture the complex interdependencies between different market factors.
8. Outlook and Future Opportunities
In this section, we point out some future research directions of diffusion models for time series and spatio-temporal data that are worthy of further investigation. (1) Scalability and Efficiency: Diffusion models remain computationally demanding, limiting their use in resource-constrained settings. Future work should focus on developing lighter architectures, enhancing sampling, model compression, and parallelization to improve real-world applicability. (2) Robustness and Adaptability: Real-world time series often contain noise, missing values, anomalies, and distribution shifts. Strengthening the robustness and generalization of diffusion models across datasets and scenarios, as well as enabling dynamic adaptation to changing environments, is essential. (3) Domain-Generalizable Challenges: Time series benchmarks are typically small and exhibit strong autocorrelation, causing overfitting. Despite the regularization effect of stochastic denoising, generalization to unseen domains remains limited. The recently identified ”accuracy law” (Wang et al., 2025b) suggests that many benchmarks may approach their intrinsic predictability limits, highlighting risks of saturation. Future work should investigate domain-generalizable diffusion frameworks via temporal augmentation, cross-domain transfer, and prior-guided regularization, advancing toward foundation models for heterogeneous temporal modalities. (4) Prior Knowledge Guided Generation: Generated time series, trajectories, and spatio-temporal patterns should satisfy domain-specific constraints (e.g., road networks, population dynamics, thermodynamics), yet current diffusion models insufficiently incorporate such prior knowledge. (5) Multimodal Data Fusion: Time series often co-occur with text or images. Integrating multimodal information can enhance analysis in applications such as finance and healthcare. Thus, new architectures are needed for effective multimodal fusion. (6) Emerging Trends and Cross-Domain Inspirations: Advances in image-based diffusion, such as latent diffusion (Rombach et al., 2021), consistency models, and rectified flows (Song et al., 2023; Armegioiu et al., 2025), offer efficiency and controllability insights, though direct transfer is hindered by temporal causality and non-stationarity. Adapting these ideas to time-aware diffusion and exploring multimodal conditioned temporal generation represent key directions. (7) Integration of LLMs and Diffusion Models: Combining LLMs with diffusion models may enhance temporal reasoning and support more comprehensive decision-making by uniting semantic understanding with generative modeling. Future work could develop hybrid architectures leveraging both capabilities.
9. Conclusion
In this survey, we presented a comprehensive overview of the advancements and applications of diffusion models in the context of time series and spatio-temporal data analysis. We categorized diffusion models into unconditioned and conditioned types, each offering distinct advantages and challenges. Furthermore, we examined the various tasks associated with these models, including forecasting, generation, imputation, and anomaly detection. Additionally, we explored different application scenarios and provided insights into future opportunities and directions in this research field. It is our hope that this survey will contribute to the advancement of research in the area of diffusion models for time series and spatio-temporal data analysis.